# On the Emergence of Linear Analogies in Word Embeddings

### v1

@@ 1. Introduction and Motivation
Key observations to explain are that (i) the top eigenvectors of $M$ already separate semantic aspects of words [6]; (ii) If the word embedding is built by approximating $M\approx W^{\top}W$, where $W$ is a $K\times m$ matrix whose columns $W_{i}$ are the embeddings of each word $i=1...m$ and rows $W^{\top}_{i}$ are the (scaled) eigenvectors of $M$, then one finds that linear analogies initially becomes more accurate as $K$ increases, such that more of the top eigenvectors of $M$ are retained [6]; (iii) linear analogies improve further when the PMI is used instead of $M$ itself, i.e. by approximating the $PMI\approx W^{\top}W$ [3, 7]. Moreover, (iv) linear analogies persists even when the word pairs exemplifying the target analogy are removed from the dataset [8] (e.g. removing from the corpus all word pairs including a masculine word and its feminine counterpart, such as man-woman, king-queen, etc.).
@@ 2. Our Contributions
(3). We find that our conclusions are remarkably robust to the presence of noise or to perturbations of the co-occurrence matrix, including the addition of i.i.d. noise, the pruning of most words, as well as the removal of co-occurrence probabilities that compose an analogy (such as “king”-“man” and “queen”-“woman”), showing how the analogy-supporting subspace remains robust.
@@ 3. Related Work
Extensive empirical evidence suggests that many models of natural language, including large language models, exhibit striking linear structure in their latent space [9, 10, 11]. This observation motivates modern research in modern language models, including mechanistic interpretability [12, 13, 14], in-context-learning [15, 16, 17], and LLM alignment [18, 19, 20]. Linear analogy structure in word embedding models is the natural precursor to these phenomena; thus, to understand linear representations in general, it is important to develop a theoretical understanding of linear analogies in simple models.
@@ 4. A model for words co-occurrence
Our main assumption is that the occurrence statistics of a word $i$ is governed by the set of attributes that define it [28], such as feminine v.s. masculine, royal v.s. non-royal, adjective v.s. noun, etc. For simplicity, we consider that there are $d$ binary attributes (extending the model to attributes with more than two choices is straightforward, as discussed in Appendix). A word is thus represented as a vector $\bm{\alpha}_{i}\in\{-1,+1\}^{d}$ of the $d$-dimensional hypercube. As a first step, we assume that all the $2^{d}$ possible words exist. We will see below that even if the vocabulary is much smaller than the total number of possible words, our conclusions hold.
We further assume that different attributes affect the statistics of words in an independent manner. As a result, the probability $P(i)$ of word $i$ follows from the set $\{p_{k}\leq 1/2,k=1...d\}$, where $p_{k}$ indicates the probability that attribute $k$ is $+1$:
$P(i)=\prod_{k=1}^{d}(p_{k}\delta_{\alpha_{i}^{(k)},1}+(1-p_{k})\delta_{\alpha_{i}^{(k)},-1})$
Likewise, the probability $P(i,j)$ that words $i$ and $j$ co-occur follows from the probability $P^{(k)}(\alpha_{i}^{(k)},\alpha_{j}^{(k)})$ that two words with given attributes $\alpha_{i}^{(k)}$ and $\alpha_{j}^{(k)}$ co-occur:
@@ 5. Word embedding directly from the co-occurrence matrix
Proof sketch: This theorem follows from the fact that the matrix $M$ is a Kronecker product of the $d$ matrices $P^{(k)}$, and from standard results on the eigen-decomposition of these products [29]. The block structure of $M$ defining a Kronecker product is most apparent if we order each word $i$ based on the lexicographic order of its attributes $\bm{\alpha}_{i}$. That is, the first coordinate that differs between $\bm{\alpha}_{i}$ and $\bm{\alpha}_{j}$ will determine which word appears first, with $i$ appearing first if its coordinate is $+1$. See appendix for details.
@@
Eigensystem of $M$: The spectrum of $M$ can be obtained using Equation 9 in Equation 8 and the corresponding eigenvectors by using Equation 10 in Equation 7. In the limit where the signals $\{s_{k},k=1...d\}$ are small, the top eigenbands are:
@@
As a result, the spectrum of $M$ depends on the distribution of the $s_{k}$. A representative $M$ with minimal variance in that distribution is shown in Figure 1b, in the symmetric case where $q_{k}=1$ for all $k$, and can be contrasted with the empirical Wikipedia derived co-occurrences reported in Figure 1a. Figure 1c shows the density of eigenvalues, averaged over 50 realizations. When the variance in the distribution in the $s_{k}$’s is small, the eigenvalue spectrum is resolved into discrete bands, while at higher values of the variance in the $s_{k}$, these bands merge and the distribution is well-fitted by a log-normal distribution.
@@ 6. Word embedding from the pointwise mutual information (PMI) matrix
Result 4: The eigenvalues of PMI are ${\cal O}(2^{d})$. It follows from the fact that $v_{k}^{T}(\log M)v_{k}={\cal O}(2^{d})$ for $k=0,...,d$. When $\bm{\eta}$ is small ($\bm{\eta}=0$ if $q_{k}=1$ for all $k$), the matrix $\log M$ has a simple spectral structure, with a top eigenvalue $\lambda_{0}=2^{d}\delta$ whose eigenvector is approximately $v_{0}$; and $d$ attribute eigenvalues $\lambda_{k}=2^{d}\gamma_{k}$ which correspond approximately to the $v_{k}$’s. All other eigenvalues are zero.
@@
Numerical validation: The inset of Figure 1c confirms that the spectrum of the PMI matrix lacks the higher-order modes present in the spectrum of $M_{ij}$. For each realization of the PMI matrix in the symmetric case simulated here, there are exactly $d$ semantically relevant eigenvalues in the spectrum.
@@
[…]
@@
In Figure 2a we report analogy accuracy using embeddings derived from Wikipedia text co-occurence matrices. The Mikolov et al. analogy task set consists of 19,544 sets of four words analogies, e.g. “hand:hands::rat:rats”, divided among 13 families e.g. adjective-superlative, verb-participle, country-nationality, etc. In Figure 2b we report the average performance for embeddings obtained with different $M_{ij}$ and $\log(M_{ij})$. The $\log(M)$ target performs strictly better than the raw $M$ target, and saturates at high embedding dimension.
@@
In Figure 2c we show the emergence of linear analogical reasoning for a single realization of this model for different fixed $k_{1}$ attributes, averaged over the $k_{2}\neq k_{1}$. The broadly distributed $s_{k}$’s give rise to analogy performances that vary considerably with embedding dimension. In Figure 2d, we show the performance for the same targets that are used in the text data. Validating our results, the $\log(M)$ target achieves 100% accuracy, regardless of the $s_{k}$ distribution for $K\geq d$. This is better than the $M_{ij}$ target, because performance there is only good when entire eigenbands of the spectrum are included (as we show in the appendix). For broadly distributed $s_{k}$ (as occurs in real text data, cf. Figure- 1), the bands mix and linear analogical reasoning is lost at increasing $K$.
@@ 7. Additive Noise Perturbation to the PMI matrix
where $\xi(i,j)$ are independent, zero-mean random variables with bounded variance $\mathbb{E}[\xi(i,j)^{2}]=\sigma_{\xi}^{2}$, and $\xi(i,j)=\xi(j,i)$ to preserve symmetry. We have: $\log M^{\prime}=\log M+\Delta$ with $\Delta$ the symmetric noise matrix. The spectral norm of the random matrix $\Delta$ is asymptotically $||\Delta||_{2}\sim 2\sigma_{\xi}\sqrt{2^{d}}$ (see theorem 2.1 of [30]). At fixed $\sigma_{\xi}$ in the limit of large $d$, this norm is negligible in comparison with distance between the semantic eigenvalues, which is ${\mathcal{O}}({2^{d}}/d)$ in the non-degenerate case. We can invoke the eigenvalue stability inequality, $|\lambda_{k}(\log M^{\prime})-\lambda_{k}(\log M)|\leq||\Delta||_{2}\sim 2^{d/2}$ (a consequence of Weyl’s inequality, see[31]), to justify that the eigen-decomposition of $M$ is thus not affected in the limit of large $d$. The linear analogies of Eq.11 is thus approximately preserved in that limit.
@@ 8. Spectral Stability under Vocabulary Subsampling
Numerical validation: in Figure- (2f), we report results on a sparsified variation of the model in $d=12$ for retention probability $f=0.15$ (see Appendix for a representative sparsified co-occurrence matrix). Despite both a multiplicative noise and the removal of $>97$% of the co-occurrence matrix, excellent analogical reasoning performance can be obtained for $K\geq d$.
@@ 9. Analogy structure after removal of all pairs of a given relationship
Numerical validation: In Figure 3a we show the performance of the PMI matrix for different analogy families. For each family, we construct a pruned PMI matrix, with co-occurrences of pairs matching the analogy set to zero, and find a minimal performance degradation in analogy accuracy in Figure 3b. We try this experiment in the model, pruning either the strongest semantic direction (WLOG, $k=1$) or the weakest semantic direction (WLOG) $k=d$ in Figure 3c. In both cases, linear analogies survive this perturbation, with perfect accuracy appearing at $K=d$. Analogies that include the pruned dimension perform similarly to analogies that do not include the pruned dimension. As with the real data (Figure 3b), this pruning introduces additional noisy eigenvectors to the representation at high embedding dimension which eventually leads to a breakdown of linear analogies.
@@ 10. Discussion
Word embeddings are central to the interpretation of large language models. Surprisingly, linear subspaces characterizing semantically meaningful concepts, originally found in classical word embedding methods, are also realized in large language models [9, 33, 34, 35] where they can enable control of model behavior [13]. The prevalence of these linear subspaces suggest that the statistics of language plays a fundamental role, and revisiting classical word embedding methods allows us to develop a sharper understanding of this question. We have shown that linear analogies in the word embeddings in such algorithms naturally arise if words are characterized by a list of attributes, and if each attribute affects the context of their associated word in an independent manner. We formulated a simple, analytically tractable model of co-occurrence statistics that captures this view, where attributes are binary and all combinations of attributes correspond to a word. This model rationalizes various observations associated with the emergence of linear analogies, and provides a fine-grained description of how they depend on the embedding dimension and the co-occurrence matrix considered. Remarkably, the model is extremely robust to perturbations including noise, the sparsification of the word vocabulary, or the removal of all pairs associated to a specific relation.
@@
The code used to produce the model results is included in the supplementary, Wikipedia co-occurrence statistics, and figures are included in the supplementary information. All simulations together run in under 150 minutes on an Nvidia H100.
@@
In the special (symmetric) case that $q=1$, the eigenvalues are simply $2$ and $2s$ for the -- and $+$ cases respectively with eigenvectors $v_{\pm}=[1,\pm 1]^{\top}$
@@
Theorem: The matrix $M(i,j)=P(i,j)/(P(i)P(j))=\prod^{d}_{k}P^{(k)}(\alpha_{i}^{(k)},\alpha_{j}^{(k)})$ indexed by word $i$ of attributes $\bm{\alpha}_{i}$ and word $j$ of attributes $\bm{\alpha}_{j}$ defined by Equation 16 has eigenvectors
@@
where the $\lambda_{\pm}^{(k)}$ are eigenvalues of the $2\times 2$ matrices $P^{(k)}$ defined in Equation 16, and $v_{\pm}^{(k)}$ are the two eigenvectors of these matrices given in Equation 18.
@@
By the definition in Equation 22, we have:
@@
Using the row-indexing notation of Equation 23, this expression can be written as:
@@
and using the indexing identity of Equation 24, this simplifies to
@@
where the attribute $s_{1}$ corresponds to “Royalty” and $s_{2}$ to “Gender”. The word embeddings for $K=2$ and $K=3$ are given in Table 1.
@@
which is satisfied uniquely when $K=3$, but not when $K=2$ as is depicted in fig. 5 and fig. 5. As the example here shows, each successive eigenvector (after the first trivial $\mathbf{1}$ vector) included in the representation allows the embedding to disambiguate one additional semantic direction – with $K=2$ only the royalty axis present in the representation, while in $K=3$, both the royalty and gender axes are present.
$w_{\text{Woman}}=w_{\text{Man}}=2\,(1,-\sqrt{s_{1}})$$O$$w_{\text{Queen}}=w_{\text{King}}=2\,(1,\sqrt{s_{1}})$$\bm{w}_{\text{Man}}-\bm{w}_{\text{Woman}}$
@@
$\bm{w}_{\text{Woman}}=2\,(1,-\sqrt{s_{1}},\sqrt{s_{2}})$$\bm{w}_{\text{Man}}=2\,(1,-\sqrt{s_{1}},-\sqrt{s_{2}})$$O$$\bm{w}_{\text{Queen}}=2\,(1,\sqrt{s_{1}},\sqrt{s_{2}})$$\bm{w}_{\text{King}}=2\,(1,\sqrt{s_{1}},\sqrt{s_{2}})$$\bm{w}_{\text{Man}}-\bm{w}_{\text{Woman}}$
@@
We now generalize our model to the case where each semantic attribute $\alpha^{(k)}$ can take one of three values: $-1$, 00, or $+1$. The value 00 corresponds to a neutral setting, indicating that the word does not express this semantic dimension. This extension allows us to model richer vocabularies where many attributes may be inactive for a given word. For simplicity of notations, we consider the case where the three values of each attribute are equally likely.
@@
Here: - $s_{k}\in[0,1]$ encodes the strength of similarity between matching vs. opposite polarities ($+1$ vs. $-1$), - $f_{k}\ll 1$ is the frequency of non-neutral settings for attribute $k$.
@@
$v_{1}$ $=\frac{1}{\sqrt{2}}(1,-1,0)^{\top}$ (contrast between $-1$ and $+1$)
@@
- The contrast direction between $-1$ and $+1$ preserves the same eigenvalue as in the binary case: $2s_{k}$,
@@
In Figure 6, we measure the analogy accuracy for narrowly distributed semantic strengths $s_{k}\sim\mathcal{N}(1/2,10^{-3})$. Accuracy with the $M_{ij}$ matrix target is perfect whenever a complete eigenband is included in the representation. Perfect analogy reconstruction is possible for the symmetric model, but is very fragile and requires a narrow distribution of $s_{k}$.
When there is greater variance in the $s_{k}$ distribution, e.g. $s_{k}\in(0,1)$, then the higher-order eigenvalues (corresponding to $s_{k}s_{k^{\prime}}$) begin to be included, without first capturing all of the first-order eigenvalues. In this case, perfect analogy accuracy is typically not attained and the peak in accuracy happens at higher $d$, as can be seen in Figure 7. In the $\log(M)$ case, the semantic eigenvalues are all included at $K=d$, as is indicated by the vertical lines in  Figure 7.
@@
In Figure 8, we show the effect of sparsification on the co-occurrence matrix $M$ at a fraction $f=0.15$. Each row $i$ in $M$ corresponds to co-occurrences including word $i$. By removing row $i$ and column $i$, we effectively remove word $i$ from the vocabulary. Despite removing $\approx 98\%$ of the co-occurrence matrix, the overall hierarchical structure is still readily apparent. We test the effect of sparsification on the top $d$ semantic eigenvalues of a $\log(M)$ target in Figure 9. As is predicted by the theory, the top $d$ eigenvalues are minimally perturbed by sparsification beyond a simple rescaling of $f=m/2^{d}$. This breaks down only when the number of retained words approaches $m\rightarrow d$, meaning that a tremendously sparsified vocabulary still retains the same eigenvalue structure.
@@
Here, we study the effect of introducing a noise perturbation $\log(M_{ij}^{\prime})=\log(M_{ij})+\xi_{ij}$ onto the PMI matrix, with $\xi_{ij}\sim\mathcal{N}(0,\sigma_{\xi})$. The spectral density has two prominant features: a set of “semantic” eigenvalues, with $\lambda_{k}\sim 2^{d}\gamma_{k}$ and a set of small, noisy eigenvalues starting at a scale set by $\sigma_{s}$, as can be seen in Figure 10a.
The $2^{d}\times 2^{d}$ dimensional $\bm{\xi}$ matrix has a spectral norm converging to $||\bm{\xi}||\sim 2\sigma_{s}\sqrt{2^{d}}$ as $d\rightarrow\infty$. This scale collapses the noise floor of $\log(M^{\prime})$ across both dimension and $\sigma_{s}$ as can be seen in Figure 10c.
@@
This is confirmed by the collapse in Figure 10c.

### v2

@@ 1. Introduction and Motivation
Key observations to explain are that (i) the top eigenvectors of $M$ already separate semantic aspects of words [6]; (ii) If the word embedding is built by approximating $M\approx W^{\top}W$, where $W$ is a $K\times m$ matrix whose columns $W_{i}$ are the embeddings of each word i=1mi=1...m and rows $W^{\top}_{i}$ are the (scaled) eigenvectors of $M$, then one finds that linear analogies initially becomes more accurate as $K$ increases, such that more of the top eigenvectors of $M$ are retained [6]; (iii) linear analogies improve further when the PMI is used instead of $M$ itself, i.e. by approximating the $PMI\approx W^{\top}W$ [3, 7]. Moreover, (iv) linear analogies persists even when the word pairs exemplifying the target analogy are removed from the dataset [8] (e.g. removing from the corpus all word pairs including a masculine word and its feminine counterpart, such as man-woman, king-queen, etc.).
@@ 2. Our Contributions
(3). We find that our conclusions are remarkably robust to the presence of noise or to perturbations of the co-occurrence matrix, including the addition of i.i.d. noise, the pruning of most words, as well as the removal of co-occurrence probabilities that compose an analogy (such as “king”-“man” and “queen”-“woman”), and correlations between attributes, showing how the analogy-supporting subspace remains robust.
@@ 3. Related Work
Extensive empirical evidence suggests that many models of natural language [2, 3], including those trained on non-english corpora [9, 10, 11], and contemporary large language models [12, 13, 14, 15], exhibit striking linear structure in their latent space. This observation motivates contemporary research in modern language models, including mechanistic interpretability [16, 17, 18], in-context-learning [19, 20, 21], and LLM alignment [22, 23, 24]. Linear analogy structure in word embedding models is the natural precursor to these phenomena; thus, to understand linear representations in general, it is important to develop a theoretical understanding of linear analogies in simple models.
@@ 4. A model for words co-occurrence
Our main assumption is that the occurrence statistics of a word $i$ is governed by the set of attributes that define it [32], such as feminine v.s. masculine, royal v.s. non-royal, adjective v.s. noun, etc. For simplicity, we consider that there are $d$ binary attributes (extending the model to attributes with more than two choices is straightforward, as discussed in Appendix A3). The word $i$ is thus represented as a vector $\bm{\alpha}_{i}\in\{-1,+1\}^{d}$ of the $d$-dimensional hypercube. As a first step, we assume that all the $2^{d}$ possible words exist. We will see below that even if the vocabulary is much smaller than the total number of possible words, our conclusions hold.
We further assume that different attributes affect the statistics of words in an independent manner (this assumption is relaxed in Appendix A5). As a result, the probability $P(i)$ of word $i$ follows from the set $\{p_{k}\leq 1/2,k=1...d\}$, where $p_{k}$ indicates the probability that attribute $k$ is +1+1:
$P(i)=\prod_{k=1}^{d}\left(p_{k}\delta(\alpha_{i}^{(k)},1)+(1-p_{k})\delta(\alpha_{i}^{(k)},-1)\right)$
where $\delta(i,j)$ denotes the Kronecker delta ($\delta(i,j)=1$ iff $i=j$ and zero otherwise), and $\alpha_{i}^{(k)}$ denotes the $k$th entry of $\bm{\alpha}_{i}$. Likewise, the probability $P(i,j)$ that words $i$ and $j$ co-occur follows from the probability $P^{(k)}(\alpha_{i}^{(k)},\alpha_{j}^{(k)})$ that two words with given attributes $\alpha_{i}^{(k)}$ and $\alpha_{j}^{(k)}$ co-occur:
@@ 5. Word embedding directly from the co-occurrence matrix
Proof sketch: This theorem follows from the fact that the matrix $M$ is a Kronecker product of the $d$ matrices $P^{(k)}$, and from standard results on the eigen-decomposition of these products [33]. The block structure of $M$ defining a Kronecker product is most apparent if we order each word $i$ based on the lexicographic order of its attributes $\bm{\alpha}_{i}$. That is, the first coordinate that differs between $\bm{\alpha}_{i}$ and $\bm{\alpha}_{j}$ will determine which word appears first, with $i$ appearing first if its coordinate is +1+1. See Appendix A1 for the detailed proof, a review of Kronecker products, and an explicit construction for the case $d=2$.
@@
Eigensystem of $M$: The spectrum of $M$ can be obtained using Equation˜9 in Equation˜8 and the corresponding eigenvectors by using Equation˜10 in Equation˜7. In the limit where the signals $\{s_{k},k=1...d\}$ are small, the top eigenbands are:
@@
As a result, the spectrum of $M$ depends on the distribution of the $s_{k}$. A representative $M$ with minimal variance in that distribution is shown in Figure˜1b, in the symmetric case where $q_{k}=1$ for all $k$, and can be contrasted with the empirical Wikipedia derived co-occurrences reported in Figure˜1a. Figure˜1c shows the density of eigenvalues, averaged over 50 realizations. When the variance in the distribution in the $s_{k}$’s is small, the eigenvalue spectrum is resolved into discrete bands, while at higher values of the variance in the $s_{k}$, these bands merge and the distribution is well-fitted by a log-normal distribution.
@@ 6. Word embedding from the pointwise mutual information (PMI) matrix
Result 4: The eigenvalues of PMI are ${\cal O}(2^{d})$. It follows from the fact that $v_{k}^{T}(\log M)v_{k}={\cal O}(2^{d})$ for k=0,,dk=0,...,d. When $\bm{\eta}$ is small ($\bm{\eta}=0$ if $q_{k}=1$ for all $k$), the matrix $\log M$ has a simple spectral structure, with a top eigenvalue $\lambda_{0}=2^{d}\delta$ whose eigenvector is approximately $v_{0}$; and $d$ attribute eigenvalues $\lambda_{k}=2^{d}\gamma_{k}$ which correspond approximately to the $v_{k}$’s. All other eigenvalues are zero.
@@
Numerical validation: The inset of Figure˜1c confirms that the spectrum of the PMI matrix lacks the higher-order modes present in the spectrum of $M_{ij}$. For each realization of the PMI matrix in the symmetric case simulated here, there are exactly $d$ semantically relevant eigenvalues in the spectrum.
@@
$\textbf{top-1 accuracy}=\frac{1}{|\mathcal{T}|}\sum_{(A,B,C,D)\in\mathcal{T}}\delta\left(W_{D},\operatorname*{argmin}_{D^{\prime}\in\mathcal{V}}(|W_{A}-W_{B}+W_{C}-W_{D^{\prime}}|_{2})\right)$ (13)
@@
In Figure˜2a we report analogy accuracy using embeddings derived from Wikipedia text co-occurence matrices. The Mikolov et al. analogy task set consists of 19,544 sets of four words analogies, e.g. “hand:hands::rat:rats”, divided among 13 families e.g. adjective-superlative, verb-participle, country-nationality, etc. In Figure 2b we report the average performance for embeddings obtained with different $M_{ij}$ and $\log(M_{ij})$. The $\log(M)$ target performs strictly better than the raw $M$ target, and saturates at high embedding dimension.
@@
In Figure˜2c we show the emergence of linear analogical reasoning for a single realization of this model for different fixed $k_{1}$ attributes, averaged over the $k_{2}\neq k_{1}$. The broadly distributed $s_{k}$’s give rise to analogy performances that vary considerably with embedding dimension. In Figure˜2d, we show the performance for the same targets that are used in the text data. Validating our results, the $\log(M)$ target achieves 100% accuracy, regardless of the $s_{k}$ distribution for $K\geq d$. This is better than the $M_{ij}$ target, because performance there is only good when entire eigenbands of the spectrum are included (as we show in the appendix). For broadly distributed $s_{k}$ (as occurs in real text data, cf. Figure- 1), the bands mix and linear analogical reasoning is lost at increasing $K$.
@@ 7. Additive Noise Perturbation to the PMI matrix
[…]
@@ 8. Spectral Stability under Vocabulary Subsampling
Numerical validation: in Figure- (2f), we report results on a sparsified variation of the model in $d=12$ for retention probability $f=0.15$ (see Appendix for a representative sparsified co-occurrence matrix). Despite both a multiplicative noise and the removal of >97>97% of the co-occurrence matrix, excellent analogical reasoning performance can be obtained for $K\geq d$.
@@ 9. Analogy structure after removal of all pairs of a given relationship
Numerical validation: In Figure˜3a we show the performance of the PMI matrix for different analogy families. For each family, we construct a pruned PMI matrix, with co-occurrences of pairs matching the analogy set to zero, and find a minimal performance degradation in analogy accuracy in Figure˜3b. We try this experiment in the model, pruning either the strongest semantic direction (WLOG, $k=1$) or the weakest semantic direction (WLOG) $k=d$ in Figure˜3c. In both cases, linear analogies survive this perturbation, with perfect accuracy appearing at $K=d$. Analogies that include the pruned dimension perform similarly to analogies that do not include the pruned dimension. As with the real data (Figure˜3b), this pruning introduces additional noisy eigenvectors to the representation at high embedding dimension which eventually leads to a breakdown of linear analogies.
@@ 10. Discussion
Word embeddings are central to the interpretation of large language models. Surprisingly, linear subspaces characterizing semantically meaningful concepts, originally found in classical word embedding methods, are also realized in large language models [12, 13, 37, 38] where they can enable control of model behavior [17] and are crucial for fact retrieval [39]. While these linear subspaces are sometimes encoded in a context-dependent manner [15], it is possible to obtain context-free LLM embeddings (as in [40]) to test our views in modern LLMs. The prevalence of these linear subspaces suggest that the statistics of language plays a fundamental role, and revisiting classical word embedding methods allows us to develop a sharper understanding of this question.
We have shown that linear analogies in the word embeddings in such algorithms naturally arise if words are characterized by a list of attributes, and if each attribute affects the context of their associated word in an independent manner. We formulated a simple, analytically tractable model of co-occurrence statistics that captures this view, where attributes are binary and all combinations of attributes correspond to a word. This model rationalizes various observations associated with the emergence of linear analogies, and provides a fine-grained description of how they depend on the embedding dimension and the co-occurrence matrix considered. Remarkably, the model is extremely robust to perturbations including noise, the sparsification of the word vocabulary, the introduction of correlations between semantic attributes, or the removal of all pairs associated to a specific relation.
@@
The code used to produce the model results, Wikipedia co-occurence statistics, and figures is available on GitHub at https://github.com/DJKorchinski/linear-analogies-word-embedding-reproduction and in the supplementary files. All simulations together run in under 150 minutes on an Nvidia H100.
Acknowledgments and Disclosure of Funding
D. J. Korchinski acknowledges financial support from the Natural Sciences and Engineering Research Council of Canada (NSERC PDF - 587940 - 2024). D. Karkada thanks Google DeepMind and BAIR for funding support. This work was supported by the Simons Foundation through the Simons Collaboration on the Physics of Learning and Neural Computation (Award ID: SFI-MPS-POL-00012574-05), PIs Bahri and Wyart.
@@
In the special (symmetric) case that $q=1$, the eigenvalues are simply $2$ and $2s$ for the - and ++ cases respectively with eigenvectors $v_{\pm}=[1,\pm 1]^{\top}$
@@
Theorem: The matrix $M(i,j)=P(i,j)/(P(i)P(j))=\prod^{d}_{k}P^{(k)}(\alpha_{i}^{(k)},\alpha_{j}^{(k)})$ indexed by word $i$ of attributes $\bm{\alpha}_{i}$ and word $j$ of attributes $\bm{\alpha}_{j}$ defined by Equation˜16 has eigenvectors
@@
where the $\lambda_{\pm}^{(k)}$ are eigenvalues of the $2\times 2$ matrices $P^{(k)}$ defined in Equation˜16, and $v_{\pm}^{(k)}$ are the two eigenvectors of these matrices given in Equation˜18.
@@
By the definition in Equation˜22, we have:
@@
Using the row-indexing notation of Equation˜23, this expression can be written as:
@@
and using the indexing identity of Equation˜24, this simplifies to
@@
where the attribute $s_{1}$ corresponds to “Royalty” and $s_{2}$ to “Gender”. The word embeddings for $K=2$ and $K=3$ are given in Table˜1.
@@
which is satisfied uniquely when $K=3$, but not when $K=2$ as is depicted in fig.˜5 and fig.˜5. As the example here shows, each successive eigenvector (after the first trivial $\mathbf{1}$ vector) included in the representation allows the embedding to disambiguate one additional semantic direction – with $K=2$ only the royalty axis present in the representation, while in $K=3$, both the royalty and gender axes are present.
[…]
@@
[…]
@@
We now generalize our model to the case where each semantic attribute $\alpha^{(k)}$ can take one of three values: 1-1, 0, or +1+1. The value 0 corresponds to a neutral setting, indicating that the word does not express this semantic dimension. This extension allows us to model richer vocabularies where many attributes may be inactive for a given word. For simplicity of notations, we consider the case where the three values of each attribute are equally likely.
@@
Here: - $s_{k}\in[0,1]$ encodes the strength of similarity between matching vs. opposite polarities (+1+1 vs. 1-1), - $f_{k}\ll 1$ is the frequency of non-neutral settings for attribute $k$.
@@
$v_{1}$ $=\frac{1}{\sqrt{2}}(1,-1,0)^{\top}$ (contrast between 1-1 and +1+1)
@@
- The contrast direction between 1-1 and +1+1 preserves the same eigenvalue as in the binary case: $2s_{k}$,
@@
In Figure˜6, we measure the analogy accuracy for narrowly distributed semantic strengths $s_{k}\sim\mathcal{N}(1/2,10^{-3})$. Accuracy with the $M_{ij}$ matrix target is perfect whenever a complete eigenband is included in the representation. Perfect analogy reconstruction is possible for the symmetric model, but is very fragile and requires a narrow distribution of $s_{k}$.
When there is greater variance in the $s_{k}$ distribution, e.g. $s_{k}\in(0,1)$, then the higher-order eigenvalues (corresponding to $s_{k}s_{k^{\prime}}$) begin to be included, without first capturing all of the first-order eigenvalues. In this case, perfect analogy accuracy is typically not attained and the peak in accuracy happens at higher $d$, as can be seen in Figure˜7. In the $\log(M)$ case, the semantic eigenvalues are all included at $K=d$, as is indicated by the vertical lines in  Figure˜7.
@@
In Figure˜8, we show the effect of sparsification on the co-occurrence matrix $M$ at a fraction $f=0.15$. Each row $i$ in $M$ corresponds to co-occurrences including word $i$. By removing row $i$ and column $i$, we effectively remove word $i$ from the vocabulary. Despite removing $\approx 98\%$ of the co-occurrence matrix, the overall hierarchical structure is still readily apparent. We test the effect of sparsification on the top $d$ semantic eigenvalues of a $\log(M)$ target in Figure˜9. As is predicted by the theory, the top $d$ eigenvalues are minimally perturbed by sparsification beyond a simple rescaling of $f=m/2^{d}$. This breaks down only when the number of retained words approaches $m\rightarrow d$, meaning that a tremendously sparsified vocabulary still retains the same eigenvalue structure.
@@
Here, we study the effect of introducing a noise perturbation $\log(M_{ij}^{\prime})=\log(M_{ij})+\xi_{ij}$ onto the PMI matrix, with $\xi_{ij}\sim\mathcal{N}(0,\sigma_{\xi})$. The spectral density has two prominant features: a set of “semantic” eigenvalues, with $\lambda_{k}\sim 2^{d}\gamma_{k}$ and a set of small, noisy eigenvalues starting at a scale set by $\sigma_{s}$, as can be seen in Figure˜10a.
The $2^{d}\times 2^{d}$ dimensional $\bm{\xi}$ matrix has a spectral norm converging to $\bm{\xi}||\sim 2\sigma_{s}\sqrt{2^{d}}$ as $d\rightarrow\infty$. This scale collapses the noise floor of $\log(M^{\prime})$ across both dimension and $\sigma_{s}$ as can be seen in Figure˜10c.
@@
This is confirmed by the collapse in Figure˜10c.